OpenStudy (anonymous):

Solve for x ln(x-2)+ln(x+2)-ln(x+1) =9

OpenStudy (anonymous):

ln(x-2)(x+2)/(x+1)=9 (x^2-4)/(x+1)=e^9. it is quadratic equation of x . x=[e^9+-square root{ e^18+4e^9+16}]/2